Given the points: A(2,6) B(-4,10) C(8,15)
Answer: A line perpendicular to AC, through A.

Answer by Photonjohn(42)

slope of AC = (15-6)/(8-2) = 9/6 = 3/2
Therefore slope of perpendicular is -2/3
through (2,6) is
y-6=-2/3(x-2)
y-6= -2/3x + 4/3
y=-2/3x + 22/3